OpenStudy (anonymous):

Let N= \[\Large \frac{ 4^{5}+4^{5}+4^{5}+4^{5} }{ 3^{5}+3^{5}+3^{3} }.\frac{ 6^{5}+6^{5}+6^{5}+6^{5}+6^{5}+6^{5} }{ 2^{5}+2^{5} }\] then \[\huge \log_2{N} \]is

OpenStudy (anonymous):

Please just give a hint only if possible , let me work it on my own

ganeshie8 (ganeshie8):

hint : \[a^n + a^n + a^n + \cdots a\text{ times} = a^n(1 + 1 + \cdots a\text{ times} ) = a^n(a) = a^{n+1}\]

OpenStudy (anonymous):